# Impact of chaotic magnetic field on mass-radius relation of rotating neutron stars

### v5

@@ Abstract
Observations reveal that magnetic fields on neutron stars (NSs) are in the range of $10^{8-15}$ G. Apart from being celestial bodies, NSs are normally rotating. In this work, we study the impact of a chaotic magnetic field on the mass-radius relation of the rotating NSs. We employ an equation of state of NSs with the nuclei in the crust and hyperons in the core. We use Hartle-Thorne formalism as an approximation of the rotating NSs. For the magnetic field ansatz, we use the one coupled to the energy density. We find that the magnetic field can decrease radius of NS. NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields. In contrast, the increment of the magnetic field can increase the compactness and deformation of rotating NSs.
@@ 1 Introduction
It has been widely understood that the evolution of the massive stars whose mass is 8--25 $M_{\odot}$ ended when neutron stars (NSs) formed. The corresponding formation is due to the stars’ gravitational collapse during Type-II, Ib, or Ic supernova explosion phenomena [1, 2]. NSs are dense, compact remnants of evolved stars, with degenerate fermions compressed by strong gravitational fields [3]. NSs are normally rotating as all celestial objects tend to naturally always rotate due to the law of angular momentum conservation [4]. Rapidly rotating NSs are observed as pulsars [5].
@@
NSs have a wide range of magnetic fields. Observations indicate that magnetic fields on NSs are at least in the range of $10^{8-15}$ G. New classes of pulsars such as Anomalous X-ray pulsars (AXPs) and Soft-Gamma Repeater pulsars (SGRs) have been identified producing vast magnetic fields. SGR is associated with remnants of a supernova, which is a young NS [6]. Highly magnetized NSs, i.e. magnetars, are typically slowly spinning NSs whose emission is thought to be powered by the decay of their large ($\geq 10^{12}$ G) magnetic fields [7]. Moreover, observations on some AXPs also indicate that AXPs surface magnetic field is around $10^{14}-10^{15}$ G [6].
@@
Konno et al. [8] calculated the deformation of NSs resulting from rotation and the presence of magnetic fields. In their analysis, a polytropic equation of state (EOS) was employed. Their findings indicate that for magnetars, the deformation induced by magnetic fields is dominant, whereas for typical pulsars, the magnetic effect is negligible. Mallick & Schramm [9] investigated mass corrections and the deformation of NSs under the influence of magnetic fields, treating the magnetic field as an anisotropy in the energy-momentum tensor (EMT). Interestingly, despite focusing on static NSs, their approach was based on the Hartle-Thorne (HT) formalism [10, 11], which is typically applied to rotating relativistic stars. In their work, magnetic perturbations in static configurations were treated analogously to rotational perturbations, following the HT approximation. They considered both very stiff and very soft EOS models in their study. Lopes & Menezes [12] explored the effects of chaotic magnetic fields on NSs. A key advantage of incorporating a chaotic magnetic field lies in the elimination of anisotropy, simplifying the mathematical formulation. They also introduced an ansatz where the magnetic field is coupled to the energy density of the NS. For their EOS models, they utilized two variations: one including hyperons and one without. However, the rotation of NSs was not considered in their analysis.
Building on previous studies, we investigate the influence of the chaotic magnetic field on the mass-radius relationship of rotating NSs. The rotational aspects are modeled using the HT formalism, while the magnetic field is described by the Lopes-Menezes ansatz. The EOS employed in this work is based on the realistic model proposed by Miyatsu et al. [13], which includes nuclei in the crust and hyperons in the core. The algorithm of the numerical simulation follows the one used in Refs. [14, 15].
@@ 2 Mathematical Formulations
To make the discussion self-contained, in Sect. 2.1, we briefly review the chaotic magnetic field of NSs, while the HT formalism is given in Sect. 2.2. It is important to note that in the mathematical formulations we use $G=c=1$
@@ 2.1 Chaotic magnetic field of Neutron Stars
It has been widely understood that the magnetic field could generate anisotropy within NSs [6, 9, 16]. This problem would make the HT formalism for rotational configuration becomes more complex (please see Refs. [1, 17]). For the magnetic field $B$ in the $z$-direction, it is well-known that the stress tensor is written in the form: $diag(\frac{B^{2}}{8\pi};\frac{B^{2}}{8\pi};-\frac{B^{2}}{8\pi})$, being non identical [12]. In Ref. [18], it is argued that the effect of a magnetic field can be described using the concept of pressure only in the case of a small-scale chaotic field. Under this condition, the pressure due to the magnetic field $p_{B}$ is shown to be consistent with field theory. Authors in Ref. [12] agreed to this argument. Now $p_{B}$ reads
@@
An ansatz of the magnetic field is proposed in Ref. [12], i.e.
@@
Here $B_{0}$ can be interpreted as the magnetic field at the center of the star, $\epsilon_{0}$ is the energy density at the center of the NS with maximum mass when the magnetic field is zero, $\gamma$ is any positive number (in this work we use $\gamma=10^{-3}$)  [^1] Values of $\gamma$ around $10^{-3}$ yield similar results, while setting $\gamma$ to much larger values (e.g., 1, 2, 3) significantly affects the outcome, likely due to changes in energy density affecting the core-crust transition., and $B_{surf}$ is the magnetic field at the surface of the NSs.
@@ 2.2 Hartle-Thorne Formalism for Rotating Relativistic Stars
In HT formalism, the metric reads [1, 10, 11, 15]
$ds^{2}$ $=$ $-e^{2\nu}dt^{2}+e^{2\lambda}dr^{2}+r^{2}e^{2\psi}(d\phi-\omega\>dt)^{2}+r^{2}e^{2\mu}d\theta^{2},$ (5)
@@
$e^{2\lambda}$ $=$ $\left[1+\frac{2}{r}\left(m_{0}+m_{2}P_{2}(\cos\theta)\right)\left(1-\frac{2m(r)}{r}\right)^{-1}\right]\left(1-\frac{2m(r)}{r}\right)^{-1}$ (7)
@@
The equilibrium equation of the relativistic bodies is given by the Tolman-Oppenheimer-Volkoff (TOV) equation [1, 10, 19], i.e.
@@
For a rotating configuration, $\bar{\omega}$ can be calculated by solving the following equation [10]:
@@
It is important to note that Eq. (10), Eq. (11), and Eq. (12) deal with static configuration. Moreover, due to the rotation, the radius of the star is deformed. To obtain the mass and radius corrections of the rotating NS, we have to employ Einstein field equation
@@
By calculating the Einstein field equation, for the ($tt$)-component of the $l=0$ order, we can obtain
$\frac{dm_{0}}{dr}$ $=$ $4\pi r^{2}\frac{d\epsilon}{dp}\left(\epsilon+p\right)\mathscr{P}_{0}+\frac{1}{12}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}-\frac{1}{3}r^{3}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}.$ (16)
From the ($rr$)-component of the $l=0$ order, we can obtain
$\frac{d\mathscr{P}_{0}}{dr}$ $=$ $-\frac{m_{0}\left(1+8\pi r^{2}p\right)}{\left(r-2m\right)^{2}}-\frac{8\pi r^{2}\left(\varepsilon+p\right)}{2\left(r-2m\right)}\mathscr{P}_{0}+\frac{1}{12}\frac{r^{4}j^{2}}{r-2m}\left(\frac{d\bar{\omega}}{dr}\right)^{2}+\frac{1}{3}\left[\frac{d}{dr}\left(\frac{r^{3}j^{2}\bar{\omega}^{2}}{r-2m}\right)\right].$ (17)
@@
where $R$ denotes the radius of the star, and $L$ denotes the angular momentum, which satisfies
$L=\frac{R^{3}}{2}\left(\Omega-\bar{\omega}(R)\right).$ (19)
@@
$M=M_{STAT}+\delta M,$ (20)
where $M_{STAT}$ is the mass of NSs within static configuration which is obtained by solving Eq. (10), Eq. (11), and Eq. (12) simultneously.
From the ($\theta\theta$), ($\phi$,$\phi$), ($\theta r$), and ($rr$) components of the $l=2$ order of Einstein field equation, we obtain
$\frac{dv_{2}}{dr}=-2\frac{d\varphi}{dr}h_{2}+\left(\frac{1}{r}+\frac{d\varphi}{dr}\right)\left[-\frac{1}{3}r^{3}\frac{dj^{2}}{dr}\bar{\omega}^{2}+\frac{1}{6}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}\right],$ (21)
$\frac{dh_{2}}{dr}$ $=$ $-2\frac{d\varphi}{dr}h_{2}+\frac{\left[8\pi r(\epsilon+p)-\frac{4m}{r^{2}}\right]}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}h_{2}-\frac{4v_{2}}{r(r-2m)\left(2\frac{d\varphi}{dr}\right)}$ (22)
$+\frac{r^{3}j^{2}}{6}\left[\frac{d\varphi}{dr}r-\frac{1}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
@@
Figure 1: Impact of chaotic magnetic field on mass-radius relation of NSs for static configuration.
(a)                                                                (b)
Figure 2: Impact of chaotic magnetic field on mass-radius relation of NSs for (a) polar radius, and (b) equatorial radius.
There exists a formula [1, 10]
@@
Finally, we can calculate the radius of the pole $R_{POL}$ and the radius of the equator $R_{EQ}$ [1, 20]

### v7

# Impact of chaotic magnetic field on physical properties of rotating neutron stars

@@ Abstract
Observations reveal that magnetic fields on neutron stars (NSs) are in the range of $10^{8-15}$ G. Apart from being celestial bodies, NSs are normally rotating. In this work, we study the impact of a chaotic magnetic field on the physical properties of the rotating NSs. i.e. mass, radius, Kepler frequency, and moment of inertia. We employ an equation of state (EOS) of NSs with the relativistic mean-field (RMF) model. Hartle-Thorne formalism as an approximation of the rotating NSs is utilized. We find that the magnetic field can decrease radius of NS. NSs formed with magnetic field exhibit a lower maximum mass compared to those without magnetic field. In contrast, the increment of the magnetic field can increase the compactness and deformation of rotating NSs. The presence of chaotic magnetic field enhances the Kepler frequency of rotating NSs, whereas it simultaneously tend to decreases their moment of inertia. The moment of inertia of rotating NSs is consistent with the constraint range obtained from pulsar mass measurements, gravitational wave event data, and X-ray observations of emissions from hotspots on NS surfaces measured by NICER.
@@ 1 Introduction
It has been widely understood that the evolution of the massive stars whose mass is 8--25 $M_{\odot}$ ends when neutron stars (NSs) form. The corresponding formation is due to the stars’ gravitational collapse during Type-II, Ib, or Ic supernova explosion phenomena Pattersons2021 ; Vidaña2018 . NSs are compact remnants of evolved stars, with degenerate fermions compressed by strong gravitational fields Reisenegger2001 . The stabilizing effect of gravity allows long-time-scale weak interactions (such as electron captures) to reach equilibrium, forming matter that is neutron-rich Graber2017 . These stars cover an extensive density range, roughly several times of nuclear density ($n_{0}\sim 10^{14}$ gm/cc), throughout their outer crust to their inner core Ghosh2024 . NSs are normally rotating as all celestial objects tend to naturally always rotate due to the law of angular momentum conservation Sivaram2012 . When rotating NSs are highly magnetized, they are recognized as pulsars Cohen1972 ; Horvath2004 ; Lyne2013 ; Basu2018 .
@@
NSs have a wide range of magnetic fields. Observations indicate that magnetic fields on NSs are at least in the range of $10^{8-15}$ G Reisenegger2001 . It is important to note that the direct observation of the magnetic field in NSs’ cores remains unachieved. However, theoretical predictions suggest that the strengths of the magnetic fields in NSs’ cores are in the order of $10^{18}$ G to $10^{20}$ G Rizaldy2018b ; Franzon2016 ; Pattersons2024b . New classes of pulsars such as Anomalous X-ray pulsars (AXPs) and Soft-Gamma Repeater pulsars (SGRs) have been identified producing vast magnetic fields. SGR is associated with remnants of a supernova, which is a young NS Rizaldy2018a . Highly magnetized NSs, i.e. magnetars, are typically slowly rotating NSs whose emission is thought to be powered by the decay of their large ($\geq 10^{12}$ G) magnetic fields, much greater than pulsars Desvignes2024 . As well as their extreme magnetic fields, magnetars are thought to have core temperature to be well above 109 K Fortin2021 . Moreover, observations on some AXPs also indicate that magnetic field of AXPs surface is around $10^{14}-10^{15}$ G Rizaldy2018a .
@@
Konno et al. Konno2000 calculated the deformation of NSs resulting from rotation and the presence of magnetic fields. In their analysis, a polytropic equation of state (EOS) was employed. Their findings indicate that for magnetars, the deformation induced by magnetic fields is dominant, whereas for typical pulsars, the magnetic effect is negligible. Mallick & Schramm Mallick2014 investigated mass corrections and the deformation of NSs under the influence of magnetic fields, treating the magnetic field as an anisotropy in the energy-momentum tensor (EMT). Interestingly, despite focusing on static NSs, their approach was based on the Hartle-Thorne (HT) formalism Hartle1967 ; Hartle1968 , which is typically applied to rotating relativistic stars. In their work, magnetic perturbations in static configurations were treated analogously to rotational perturbations, following the HT approximation. They considered both very stiff and very soft EOS models in their study. Lopes & Menezes Lopes2015 explored the effects of chaotic magnetic fields on NSs. A key advantage of incorporating a chaotic magnetic field lies in the elimination of anisotropy, simplifying the mathematical formulation. They also introduced an ansatz where the magnetic field is coupled to the energy density of the NS. For their EOS models, they utilized two variations: one including hyperons and one without hyperons. However, the rotation of NSs was not considered in their analysis.
On the other hand, in Newtonian theory, there exists a maximum rotational frequency for stars composed of perfect fluid, where gravitational and centrifugal forces are in equilibrium at their equators. This maximum frequency is proportional to the square root of the ratio of their masses to their radii Wen2007 . In the context of general relativity (GR), the angular velocity of NSs is limited by the Kepler frequency, a threshold determined by their mass and radius in static configurations Ordaz2019 ; Lopes2024 . Notably, Refs. Wen2007 ; Ordaz2019 ; Lopes2024 provide specific calculations of the Kepler frequency.
Another significant property of NSs is their moment of inertia, which offers insights into their internal structure and EOS. Given the link between the EOS and moment of inertia, approximate methods for estimating it in static models, particularly for stiff EOS scenarios, are feasible daSilva2021 . Furthermore, measuring the moment of inertia is crucial due to its universal correlation with compactness Rahmansyah2020 . Rahmansyah et al. Rahmansyah2020 calculated the moment of inertia of anisotropic NS which satisfies the constraints proposed by Landry et al. Landry2020 . However, the magnetic field is not taken into account in their calculation of the moment of inertia.
Building on previous studies, we investigate the influence of the chaotic magnetic field on some physical properties of rotating NSs, i.e. mass, radius, Kepler frequency, and moment of inertia. In particular, for the moment of inertia, we adopt the very slow rotation approximation with angular velocity $\Omega=50$ s-1. Under this condition, stellar deformation is neglected, as the stars remain nearly perfectly spherical, with eccentricity values approaching zero. This justifies the use of the standard moment of inertia formula for a spherical body. Such an approach has also been adopted in Refs. Rahmansyah2020 ; Rizaldy2024 . The rotational aspects are modeled using the HT formalism, while the magnetic field is described by the Lopes-Menezes ansatz.
In general, NS has three parts with different compositions and density ranges i.e., outer crust, inner crust and the core Sulaksono2015 . These compositions are described by an EOS. Here we employ the relativistic mean-field (RMF) EOS with BSP parameter sets Agrawal2012 . The algorithm of the numerical simulation follows the one used in Refs. Ordaz2019 ; Pattersons2024 .
@@ 2 Mathematical Formulations
To make the discussion self-contained, in Sect. 2.1, we briefly present the nuclear model used in this work; then in Sec 2.2, we shortly review the chaotic magnetic field of NSs; while the HT formalism is given in Sect. 2.2. It is important to note that in the mathematical formulations we use $G=c=1$.
2.1 Nuclear Model
The Lagrangian density for the RMF model reads Agrawal2012
$\mathcal{L}$ $=$ $\mathcal{L}_{\mathcal{N}M}+\mathcal{L}_{\sigma}+\mathcal{L}_{\omega}+\mathcal{L}_{\rho}+\mathcal{L}_{\sigma\omega\rho},$ (1)
where $\mathcal{L}_{\mathcal{N}M}$ is the Lagrangian that describes the interactions of the nucleons through the mesons. It is given by
$\mathcal{L}_{\mathcal{N}M}$ $=$ $\sum_{J=n,p}\bar{\Psi}_{J}\left[i\gamma^{\mu}\partial_{\mu}-(M-g_{\sigma}\sigma)\right.$ (2)
$\left.-\left(g_{\omega}\gamma^{\mu}\omega_{\mu}+\frac{1}{2}g_{\rho}\gamma^{\mu}\tau\cdot\rho_{\mu}\right)\right]\Psi_{J}.$
Note that the sum is taken over the neutrons and protons. Here $\tau$ are the isospin matrices. Moreover, the Lagrangian that describes the self-interactions for $\sigma$, $\omega$, and $\rho$ mesons can be written as,
$\mathcal{L}_{\sigma}$ $=$ $\frac{1}{2}(\partial_{\mu}\sigma\partial^{\mu}\sigma-m_{\sigma}^{2}\sigma^{2})-\frac{\kappa_{3}}{6M}g_{\sigma}m_{\sigma}^{2}\sigma^{3}$ (3)
$-\frac{\kappa_{4}}{24M^{2}}g_{\sigma}^{2}m_{\sigma}^{2}\sigma^{4},$
$\mathcal{L}_{\omega}$ $=$ $-\frac{1}{4}\omega_{\mu\nu}\omega^{\mu\nu}+\frac{1}{2}m_{\omega}^{2}\omega_{\mu}\omega^{\mu}$ (4)
$+\frac{1}{24}\zeta_{0}g_{\omega}^{2}(\omega_{\mu}\omega^{\mu})^{2},$
$\mathcal{L}_{\rho}$ $=$ $-\frac{1}{4}\rho_{\mu\nu}\cdot\rho^{\mu\nu}+\frac{1}{2}m_{\rho}^{2}\rho_{\mu}\cdot\rho^{\mu},$ (5)
Here, $\omega_{\mu\nu}$ and $\rho^{\mu\nu}$ are the field tensors corresponding to the $\omega$ and $\rho$ mesons, respectively. They are defined as $\omega_{\mu\nu}=\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu}$ and $\rho^{\mu\nu}=\partial^{\mu}\rho^{\nu}-\partial^{\nu}\rho^{\mu}$. Furthermore, the cross interactions of $\sigma$, $\omega$, and $\rho$ mesons are given by $\mathcal{L}_{\sigma\omega\rho}$, i.e.
$\mathcal{L}_{\sigma\omega\rho}$ $=$ $\frac{\eta_{1}}{2M}g_{\sigma}m_{\omega}^{2}\sigma\omega_{\mu}\omega^{\mu}+\frac{\eta_{2}}{4M^{2}}g_{\sigma}^{2}m_{\omega}^{2}\sigma^{2}\omega_{\mu}\omega^{\mu}$ (6)
$+\frac{\eta_{\rho}}{2M}g_{\sigma}m_{\rho}^{2}\sigma\,\rho_{\mu}\cdot\rho^{\mu}$
$+\frac{\eta_{1\rho}}{4M^{2}}g_{\sigma}^{2}m_{\rho}^{2}\sigma^{2}\,\rho_{\mu}\cdot\rho^{\mu}$
$+\frac{\eta_{2\rho}}{4M^{2}}g_{\omega}^{2}m_{\rho}^{2}\omega_{\mu}\omega^{\mu}\,\rho_{\mu}\cdot\rho^{\mu}.$
Table 1: BSP parameter sets Agrawal2012 for the RMF model.
Value
$g_{\sigma}/4\pi$ 0.8764
$g_{\omega}/4\pi$ 1.1481
$g_{\rho}/4\pi$ 1.0508
$\kappa_{3}$ 1.0681
$\kappa_{4}$ 14.9857
$\eta_{1}$ 0.0872
$\eta_{2}$ 3.1265
$\eta_{\rho}$ 0.0
$\eta_{1\rho}$ 0.0
$\eta_{2\rho}$ 53.7642
$\zeta_{0}$ 0.0
$m_{\sigma}/M$ 0.5383
$m_{\omega}/M$ 0.8333
$m_{\rho}/M$ 0.8200
The values of the free parameters of the Lagrangian for this model are listed in Table 1, while the predicted properties at saturation density—namely the binding energy per baryon ($B/A$), incompressibility ($K$), symmetry energy ($S_{0}$), the slope of the symmetry energy ($L$), and saturation density ($n_{0}$) itself—are summarized in Table 2. The constraints for the predicted values at saturation density are taken from Refs. Lopes2024 ; Rahmansyah2020 ; Dutra2014 .
Table 2: Predictions for the RMF EOS with BSP parameter sets at saturation density. the constraints are taken from Refs. Lopes2024 ; Rahmansyah2020 ; Dutra2014 .
Physical Quantities Constraints This Model
$n_{0}$ (fm-3) 0.148–0.170 0.149
$K$ (MeV) 220–260 230.97
$S_{0}$ (MeV) 28.5–34.9 28.86
$B/A$ (MeV) 15.8–16.5 15.90
$L$ (MeV) 44–116 50.00
For the crust of NSs, the one in Refs Mart2013 ; Suparti2017 is used. At the core-crust transition, the pressures of the core and the inner crust are matched. The core-crust transition occurs at the transition nuclear density $n_{t}=$ 0.087 fm-3 and transition pressure $p_{t}=$ 0.29 MeV fm-3 Rahmansyah2020 .
It is important to note that the effect of the magnetic field on the EOS is not taken into account, as our focus is solely on investigating the impact of the chaotic magnetic field on general relativistic aspects of NSs and their implication on the NSs’ physical properties.
@@ 2.2 Chaotic Magnetic Field of Neutron Stars
Generally, chaotically tangled magnetic field lines are thought to be present throughout astrophysical plasmas: planets, stars, accretion discs, galaxies, clusters of galaxies, and the intergalactic medium Kumar2017 . In nature, chaotic magnetic fields could be generated by asymmetric current configuration Dasgupta2014 .
It has been widely understood that the magnetic field could generate anisotropy within NSs Rizaldy2018b ; Rizaldy2018a ; Mallick2014 . This problem would make the HT formalism for rotational configuration becomes more complex (please see Refs. Pattersons2021 ; Posada2024 ; Becerra2024 ). For the magnetic field $B$ in the $z$-direction, it is well-known that the stress tensor is written in the form: $diag(\frac{B^{2}}{8\pi};\frac{B^{2}}{8\pi};-\frac{B^{2}}{8\pi})$, being non identical Lopes2015 . In Ref. Zeldovich , it is argued that the effect of a magnetic field can be described using the concept of pressure only in the case of a small-scale chaotic field. Under this condition, the pressure due to the magnetic field $p_{B}$ is shown to be consistent with field theory. Ref. Lopes2015 agreed to this argument. Now $p_{B}$ reads
@@
An ansatz of the magnetic field is proposed in Ref. Lopes2015 , i.e.
@@
Here $B_{0}$ can be interpreted as the magnetic field at the center of the star, $\epsilon_{0}$ is the energy density at the center of the NS with maximum mass when the magnetic field is zero, $\gamma$ is any positive number (in this work we use $\gamma=$ 3, 5, and 6), and $B_{surf}$ is the magnetic field at the surface of the NSs.
@@ 2.3 Hartle-Thorne Formalism for Rotating Relativistic Stars
The details of the HT formalism can be found in the original papers Hartle1967 ; Hartle1968 . Other interesting additional discussions on the formalism can be referred to Refs. Pattersons2021 ; Pattersons2024 ; Posada2024 ; Becerra2024 . Here we present only a little discussion and the necessary equations to calculate some physical properties, i.e. mass, radius, Kepler frequency, and moment of inertia. This is especially useful to help the readers who are not familiar with the HT formalism.
It has to be noted that based on Eq. (8) and Eq. (9), it can be seen that the total energy density and total pressure arising from the matter and the presence of a magnetic field can be redefined by introducing $\epsilon$ and $p$. Therefore, in this subsection, the definitions of $\epsilon$ and $p$ also encompass the total contributions from both matter and the magnetic field. Consequently, the formulation of HT presented in this subsection is indirectly modified by the presence of the magnetic field. Although this modification is not explicitly apparent, the contribution of the magnetic field is embedded within $\epsilon$ and $p$.
In HT formalism, the metric reads Pattersons2021 ; Hartle1967 ; Hartle1968 ; Pattersons2024
$ds^{2}$ $=$ $-e^{2\nu}dt^{2}+e^{2\lambda}dr^{2}+e^{2\psi}(d\phi-\omega\>dt)^{2}$ (11)
$+e^{2\mu}d\theta^{2},$
@@
$e^{2\lambda}$ $=$ $\left[1+\frac{2}{r}\left(m_{0}+m_{2}P_{2}(\cos\theta)\right)\left(1-\frac{2m(r)}{r}\right)^{-1}\right]$ (13)
×$\times\left(1-\frac{2m(r)}{r}\right)^{-1}$
@@
The equilibrium equation of the relativistic bodies is given by the Tolman-Oppenheimer-Volkoff (TOV) equation Pattersons2021 ; Hartle1967 ; Oppenheimer1939 , i.e.
@@
For a rotating configuration, $\bar{\omega}$ can be calculated by solving the following equation Hartle1967 :
@@
Figure 1: Mass-radius relation of NSs for static configuration with different parameter sets of chaotic magnetic field strengths.
(a)                                                   (b)
Refer to caption Refer to caption
                    (c)                                                   (d)
Refer to caption Refer to caption
                    (e)                                                   (f)
Figure 2: Mass-radius relation of NSs for (a) polar radius at $\Omega=$ 1000 s-1, (b) equatorial radius at $\Omega=$ 1000 s-1, (c) polar radius at $\Omega=$ 2000 s-1, (d) equatorial radius at $\Omega=$ 2000 s-1, (e) polar radius at $\Omega=$ 3000 s-1, and (f) equatorial radius at $\Omega=$ 3000 s-1.
It is important to note that Eqs. (16)–(18) deal with static configuration. Moreover, due to the rotation, the radius of the star is deformed. To obtain the mass and radius corrections of the rotating NS, we have to employ Einstein field equation
@@
By calculating the Einstein field equation, for the ($tt$) component of the $l=0$ order, we can obtain
$\frac{dm_{0}}{dr}$ $=$ $4\pi r^{2}\frac{d\epsilon}{dp}\left(\epsilon+p\right)\mathscr{P}_{0}+\frac{1}{12}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}$ (22)
$-\frac{1}{3}r^{3}\left(\frac{dj^{2}}{dr}\right)\bar{\omega}^{2}.$
From the ($rr$) component of the $l=0$ order, we can obtain
$\frac{d\mathscr{P}_{0}}{dr}$ $=$ $-\frac{m_{0}\left(1+8\pi r^{2}p\right)}{\left(r-2m\right)^{2}}-\frac{8\pi r^{2}\left(\varepsilon+p\right)}{2\left(r-2m\right)}\mathscr{P}_{0}+\frac{1}{12}$ (23)
×$\times\frac{r^{4}j^{2}}{r-2m}\left(\frac{d\bar{\omega}}{dr}\right)^{2}+\frac{1}{3}\left[\frac{d}{dr}\left(\frac{r^{3}j^{2}\bar{\omega}^{2}}{r-2m}\right)\right].$
@@
where $R$ denotes the radius of the star, and $L$ denotes the angular momentum. In Ref. Hartle1968 , there is a relation
${\Omega=\bar{\omega}(R)+\frac{2L}{R^{3}},}$ (25)
so we can obtain
$L=\frac{R^{3}}{2}(\Omega-\bar{\omega}(R)).$ (26)
Alternatively, angular momentum can be calculated as Glendenning1998
[…]
@@
$M=M_{0}+\delta M,$ (28)
where $M_{0}$ is the mass of NSs within static configuration which is obtained by solving Eqs. (16)–(18) simultaneously.
From the ($\theta\theta$), ($\phi\phi$), ($\theta r$), and ($rr$) components of the $l=2$ order of Einstein field equation, we obtain
$\frac{dv_{2}}{dr}$ $=$ $-2\frac{d\varphi}{dr}h_{2}+\left(\frac{1}{r}+\frac{d\varphi}{dr}\right)$ (29)
×$\times\left[-\frac{1}{3}r^{3}\frac{dj^{2}}{dr}\bar{\omega}^{2}+\frac{1}{6}j^{2}r^{4}\left(\frac{d\bar{\omega}}{dr}\right)^{2}\right],$
$\frac{dh_{2}}{dr}$ $=$ $-2\frac{d\varphi}{dr}h_{2}+\frac{\left[8\pi r(\epsilon+p)-\frac{4m}{r^{2}}\right]}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}h_{2}$ (30)
$-\frac{4v_{2}}{r(r-2m)}\left(2\frac{d\varphi}{dr}\right)^{-1}+\frac{r^{3}j^{2}}{6}$
×$\times\left[\frac{d\varphi}{dr}r-\frac{1}{(r-2m)\left(2\frac{d\varphi}{dr}\right)}\right]\left(\frac{d\bar{\omega}}{dr}\right)^{2}$
@@
There exists a formula Pattersons2021 ; Hartle1967
@@
Finally, we can calculate the radius of the pole $R_{POL}$ and the radius of the equator $R_{EQ}$ Pattersons2021 ; Lopes2024
@@
This allows us to calculate the eccentricity
$e=\sqrt{1-\frac{R_{POL}^{2}}{R_{EQ}^{2}}}.$ (36)
Furthermore, in the original papers of HT formalism Hartle1967 ; Hartle1968 , the maximum angular velocity of a rotating body within Newtonian gravity is mentioned, i.e.
$\Omega_{max}=\sqrt{M_{0}/R^{3}}.$ (37)
However, in the context of GR, the maximum angular velocity of NSs (which is called as Kepler frequency) is approximated by Lopes2024
$\Omega_{K}=0.65\sqrt{M_{0}/R^{3}}.$ (38)
Moreover, another important physical quantity to consider is the moment of inertia. With the angular momentum $L$ and the angular velocity relative to the distant observers $\Omega$ in our hands, we can determine the moment of inertia
$I=\frac{L}{\Omega}.$ (39)